OpenStudy (anonymous):

2x+1= square root 3x^2+5x+3 solve the equation for x. if necessary round your answer to two decimal places.

OpenStudy (anonymous):

\[2x+1=\sqrt{3x ^{2}+5x+3}\]

OpenStudy (anonymous):

i got x=2,-1

OpenStudy (anonymous):

try plugging both back into the original equation to see if they match on both sides

OpenStudy (anonymous):

just 2 does.